# Models of interacting bosons with exact ground states: a unified approach

### v1

@@ Abstract
We define an infinite class of “frustration-free” interacting lattice quantum Hamiltonians for number-conserving bosons, constructed such that their exact ground states have a density distribution specified by the Boltzmann weight of a corresponding classical lattice gas problem. By appropriately choosing the classical weights, we obtain boson representations of various known solvable models, including quantum dimer models and certain Levin-Wen string-net models, and we also systematically construct solvable models with other interesting ground states, including “quantum spin liquids,” supersolids, “Bose-Einstein insulators,” Bose liquids with “Bose surfaces”, and Bose-Einstein condensates that permit adiabatic evolution from a non-interacting limit to a Gutzwiller-projected limit.
@@
Frustration-free Hamiltonians are an important class of solvable models whose ground states simultaneously minimize all the elementary components. In the modern study of quantum many-body physics, such models play an important role. For example, they were used to prove the existence and illuminate the physical properties of various strongly correlated phases of matter Arovas and Girvin (1992), such as Affleck-Kennedy-Lieb-Tasaki type models Affleck et al. (1987); Arovas et al. (1988) of quantum disordered antiferromagnets, the Haldane pseudo-potential model Haldane (1983) for fractional quantum Hall effects, the Rokshar-Kivelson quantum dimer model Rokhsar and Kivelson (1988); Kivelson et al. (1987) for resonating valence bond states, Kitaev models Kitaev (2006) and Levin-Wen string-net models Levin and Wen (2005); Lin et al. (2021) for topological orders, Haah’s code Haah (2011) and X-cube model Vijay et al. (2016) for fractons, “quantum geometric nesting” models Han et al. (2024) for fermion bilinear orders in flat-band systems, among many others Kane et al. (1991); Levin et al. (2011); de Beaudrap et al. (2010); Yang (2020); Zhang and Røising (2023); Mielke and Tasaki (1993); Lin and Levin (2015); Ware et al. (2016); Zhang et al. (2023); Schossler et al. (2024); Yao and Kivelson (2012); Misguich et al. (2002); Han and Chen (2023, 2022); Katsura et al. (2021); Koch-Janusz et al. (2013); Lin and Levin (2014); Wouters et al. (2021); Schuch et al. (2011); Tovmasyan et al. (2016); Cheng et al. (2017); Qi et al. (2015); Burnell et al. (2014); Watanabe et al. (2023). Frustration-free points also appear in the parameter space of certain strongly correlated problems, providing convenient handles for investigations of nearby regimes Ran et al. (2023, 2024a, 2024b); Yan et al. (2022a, b); Huse et al. (2003); Fradkin et al. (2004); Vishwanath et al. (2004); Samajdar et al. (2021); Verresen et al. (2021); Verresen and Vishwanath (2022); Kang and Vafek (2019); Herzog-Arbeitman et al. (2022); Bultinck et al. (2020); Mao et al. (2023); Han and Kivelson (2023, 2024). For example, the prediction of a quantum spin liquid in a Rydberg array Semeghini et al. (2021) was made using this sort of methodology Verresen et al. (2021); Samajdar et al. (2021); Verresen and Vishwanath (2022). In recent years, the special entanglement structure of such models Ji et al. (2011); Chen et al. (2012); Anshu et al. (2022); Bravyi and Gosset (2015) has also drawn attention to their possible applications to quantum information processing Darmawan et al. (2012); Darmawan and Bartlett (2014); Thibodeau and Clark (2023); Zhu et al. (2024); Anshu et al. (2020).
@@
However, frustration-free models are hard to find and even to diagnose (known as the quantum satisfiability problem Sattath et al. (2016); Bravyi (2006)). Systematic construction schemes of such models are thus desirable. In this work, we construct an infinite class of frustration-free, interacting lattice models for number-conserving bosons by (roughly speaking) relating the problem to Bose-Einstein condensation (BEC) of free bosons, adopting a method spiritually similar to Witten’s conjugation argument Wouters et al. (2021); Witten (1982) for generalizing supersymmetric models. The models are relatively simple and physical, consisting of boson hopping terms and short-range few-body density interactions. The resulting ground state is a BEC state with a re-weighted density distribution identical to a classical lattice gas problem. We will show that this scheme unifies many known frustration-free models, leads to their natural generalizations, and furthermore allows constructions for other interesting phases of matter. We present the key results in the main text and defer details of the derivations to the Supplemental Materials (SM).
@@
Specifically, the solvable Hamiltonian is designed by working backward from an arbitrary real function, $K(\{n\})$, of the boson occupation configuration $\{n\}$. Ultimately, we will find that $e^{-K(\{n\})}$ is the Boltzmann weight of an associated classical lattice gas problem defined on the same spatial lattice, which has the same density correlations as the ground state of the quantum problem. We identify clusters with $L$-tuples of site indices, $c=(i_{1}\dots i_{L})$ (where a site index might occur multiple times in this unordered list). The kinetic terms associated with the pair of clusters $c$ and $c^{\prime}$ is
@@
it follows that  [^1] This form is similar to that of the interacting boson models in the continuum studied in Ref. Kane et al. (1991) where the engineered Hamiltonian can be written into the form $\hat{\mathcal{H}}=$ $\hat{P}^{-1}\left[\nabla\psi^{\dagger}\right]\hat{P}\cdot\hat{P}\left[\nabla\psi\right]\hat{P}^{-1}.$ (8) with $\hat{P}$ related to a corresponding classical Boltzmann weight determined by the boson configuration.
$\hat{H}_{c,c^{\prime}}=$ $\hat{P}^{-1}\left[\hat{B}^{\dagger}_{c}-\hat{B}^{\dagger}_{c^{\prime}}\right]\hat{P}\hat{P}_{c}\hat{P}_{c^{\prime}}\hat{P}\left[\hat{B}_{c}-\hat{B}_{c^{\prime}}\right]\hat{P}^{-1}.$ (9)
Since these Hamiltonians can be written into the form of $\hat{O}^{\dagger}\hat{O}$, they are always positive semi-definite.
@@
$\left[\hat{B}_{c}-\hat{B}_{c^{\prime}}\right]\left(\hat{b}^{\dagger}_{0}\right)^{M}|0\rangle=0$ (10)
@@
is the ground-state of $\hat{H}$, since $\hat{H}$ is non-negative and $\hat{H}|M\rangle=0$. We have not proven that this is the only ground state within the $M$-boson sector, and indeed one can construct artificial examples in which there are multiple zero energy states, but we conjecture that for generic forms of $\tau_{c,c^{\prime}}$ and $K(\{n\})$ up to degeneracies associated with spontaneously broken symmetries or topological order - the ground-state is typically unique.
@@
Dynamical density fluctuations: Besides the ground states, the dynamics of the density fluctuations can also be inferred. For concreteness, in this and the next section, we will consider the case of translationally invariant and inversion symmetric systems on regular lattices so that the excitations can be assigned a Bloch wave vector, $\bm{q}$.
@@
$f(\bm{q})\equiv$ $\frac{1}{2\mathsf{V}}\left\langle\alpha\left|[\hat{\rho}_{\bm{q}}^{\dagger},[\hat{H},\hat{\rho}_{\bm{q}}]]\right|\alpha\right\rangle$ (21)
@@
Quantum dimer models : Gutzwiller projection is not the only case in which the present approach can be used to study model problems defined in constrained Hilbert places. Here we show how this scheme can also unify quantum dimer models (QDM). In this case, the sites are taken to lie on the edges of a given lattice (forming its generalized ‘Lieb lattice’), and the bosons are considered ‘dimers.’ The dimer covering constraint can be achieved by
@@
where $N_{\mathsf{v}}\equiv\sum_{i\in\partial\mathsf{v}}n_{i}$ is the number of dimers occupying the edges connected to the vertex $\mathsf{v}$ (represented by $i\in\partial\mathsf{v}$), and $\bar{N}=\bar{n}\mathsf{C}/2$ is the average number of $N_{\mathsf{v}}$ ($\mathsf{C}$ is the coordination number of the lattice). For $10<\bar{n}<$ 1, the corresponding $\hat{P}$ projects onto states with at most one dimer on each edge as $u\rightarrow\infty$, and either $\lfloor\bar{N}\rfloor$ or $\lfloor\bar{N}\rfloor+1$ dimers touching each vertex as $u^{\prime}\to\infty$. Specifically, for $\bar{N}=1$, a hard-core dimer constraint is enforced, and hence the reference classical statistical mechanics model is the much-studied hard-core dimer model. On the other hand, for $\bar{N}=2$, the corresponding classical model sums over non-intersecting, undirected, fully-packed loops, another famous model in classical statistical mechanics  [^5] More generally, for any $\bar{N}\leq 2$, the reference classical statistical mechanics model corresponding to undirected, non-intersecting loops at general packing can be obtained by substituting $\left(N_{\mathsf{v}}-\bar{N}\right)^{2}\to N_{\mathsf{v}}\left(N_{\mathsf{v}}-2\right)^{2}$ in Eq. 29, and then taking the familiar limit $u$ and $u^{\prime}\to\infty$..
@@
Levin-Wen string-net Models: Further extending this methodology, one can also obtain ad hoc boson representations of the Levin-Wen string-net models Levin and Wen (2005); Lin et al. (2021) (including Kitaev’s toric code Kitaev (2006)) or quantum vertex models Balasubramanian et al. (2022b, a). The strategy is to introduce several flavors of bosons on lattice edge to represent the possible ‘string types’ and then use an appropriate $K(\{n\})$ to impose a single-occupancy constraint on each edge and a constraint at each vertex so that surrounding string types obey the ‘branching rules.’ By considering the quantum resonance of the strings consistent with the rules, one can, in this way, naturally obtain the solvable Hamiltonians describing certain deconfined gauge theories.
@@
The addition of a density interaction $w(\delta n)^{2}/2$ to the effective action makes the system unstable to phase separation (as usual) for attractive $w<0$. In contrast, for repulsive $w>0$, it keeps the gapless manifold $\mathcal{M}$ but changes the boson excitation dispersion near $\mathcal{M}$ from quadratic to linear: $E(\bm{q})\sim\sqrt{wf(\bm{q})}$, which results in a spatially decaying boson Green’s function $\langle\mathrm{e}^{\mathrm{i}[\theta(\bm{r})-\theta(\bm{0})]}\rangle$ with a similar character as for a Fermi liquid with the same manifold $\mathcal{M}$ of gapless modes (e.g. a surface in 3D or a nodal line in 2D)  [^6] It should be noted that the state constructed here is different from the Bose liquid studied in Ref. Motrunich and Fisher (2007) in many aspects, including its specific heat and the relation between the particle density and the volume enclosed by the Bose surfaces. .
Discussion: In this work, we have developed a unified approach for constructing “frustration-free” Hamiltonians for number-conserving bosons whose exact ground-state properties can be inferred from the properties of an associated classical statistical mechanical problem on the same lattice. For the known solvable models, our scheme provides not only an alternative representation that might be realizable on engineered platforms like optical lattices but also a natural route from a strongly constrained problem to a free (trivial) limit of the boson problem. When the constrained limit of the problem exhibits an exotic quantum phase of matter, this evolution must drive exotic transitions out of them, whose properties may be studied with these engineered models.
@@
Acknowledgements: We thank Jonah Herzog-Arbeitman, Steven Girvin, and Vladimir Calvera for helpful communications. This work was supported in part by the U. S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, under Award No. DE-SC0020045.

### v2

@@ Abstract
We define an infinite class of “frustration-free” interacting lattice quantum Hamiltonians for number-conserving bosons, constructed such that their exact ground states have a density distribution specified by the Boltzmann weight of a corresponding classical lattice gas problem. By appropriately choosing the classical weights, we obtain boson representations of various known solvable models, including quantum dimer and vertex models, toric code, and certain Levin-Wen string-net models. We also systematically construct solvable models with other interesting ground states, including “quantum spin liquids,” supersolids, “Bose-Einstein insulators,” Bose liquids with “Bose surfaces”, and Bose-Einstein condensates that permit adiabatic evolution from a non-interacting limit to a Gutzwiller-projected limit.
@@
Frustration-free Hamiltonians are an important class of solvable models whose ground states simultaneously minimize all the elementary components. In the modern study of quantum many-body physics, such models play an important role. For example, they were used to prove the existence and illuminate the physical properties of various strongly correlated phases of matter Arovas and Girvin (1992), such as Affleck-Kennedy-Lieb-Tasaki type models Affleck et al. (1987); Arovas et al. (1988) of quantum disordered antiferromagnets, the Haldane pseudo-potential model Haldane (1983) for fractional quantum Hall effects, the Rokshar-Kivelson quantum dimer model Rokhsar and Kivelson (1988); Kivelson et al. (1987) for resonating valence bond states, Kitaev models Kitaev (2006) and Levin-Wen string-net models Levin and Wen (2005); Lin et al. (2021) for topological orders, Haah’s code Haah (2011) and X-cube model Vijay et al. (2016) for fractons, “quantum geometric nesting” models Han et al. (2024) for fermion bilinear orders in flat-band systems, among many others Ardonne et al. (2004); Kane et al. (1991); Papanikolaou et al. (2007); Castelnovo et al. (2005); Levin et al. (2011); de Beaudrap et al. (2010); Yang (2020); Zhang and Røising (2023); Mielke and Tasaki (1993); Lin and Levin (2015); Ware et al. (2016); Zhang et al. (2023); Schossler et al. (2024); Yao and Kivelson (2012); Misguich et al. (2002); Han and Chen (2023, 2022); Katsura et al. (2021); Koch-Janusz et al. (2013); Lin and Levin (2014); Wouters et al. (2021); Schuch et al. (2011); Tovmasyan et al. (2016); Cheng et al. (2017); Qi et al. (2015); Burnell et al. (2014); Watanabe et al. (2023). Frustration-free points also appear in the parameter space of certain strongly correlated problems, providing convenient handles for investigations of nearby regimes Ran et al. (2023, 2024a, 2024b); Yan et al. (2022a, b); Huse et al. (2003); Fradkin et al. (2004); Vishwanath et al. (2004); Samajdar et al. (2021); Verresen et al. (2021); Verresen and Vishwanath (2022); Kang and Vafek (2019); Herzog-Arbeitman et al. (2022); Bultinck et al. (2020); Mao et al. (2023); Han and Kivelson (2023, 2024). For example, the prediction of a quantum spin liquid in a Rydberg array Semeghini et al. (2021) was made using this sort of methodology Verresen et al. (2021); Samajdar et al. (2021); Verresen and Vishwanath (2022). In recent years, the special entanglement structure of such models Ji et al. (2011); Chen et al. (2012); Anshu et al. (2022); Bravyi and Gosset (2015) has also drawn attention to their possible applications to quantum information processing Darmawan et al. (2012); Darmawan and Bartlett (2014); Thibodeau and Clark (2023); Zhu et al. (2024); Anshu et al. (2020).
@@
However, frustration-free models are hard to find and even to diagnose (known as the quantum satisfiability problem Sattath et al. (2016); Bravyi (2006)). Systematic construction schemes of such models are thus desirable. In this work, we construct an infinite class of frustration-free, interacting lattice models for number-conserving bosons by (roughly speaking) relating the problem to Bose-Einstein condensation (BEC) of free bosons, adopting a method spiritually similar to Witten’s conjugation argument Wouters et al. (2021); Witten (1982) for generalizing supersymmetric models. The models can be relatively simple and physical, consisting of short-range boson hopping terms and few-body density interactions. The resulting ground state is a BEC state with a re-weighted density distribution identical to a classical lattice gas problem. We will show that this scheme unifies many known frustration-free models, leads to their natural generalizations, and furthermore allows constructions for other interesting phases of matter. We present the key results in the main text and defer details of the derivations to the Supplemental Materials (SM).
@@
Specifically, the solvable Hamiltonian is designed by working backward from an arbitrary real function, $K(\{n\})$, of the boson occupation configuration $\{n\}$. Ultimately, we will find that $e^{-K(\{n\})}$ is the Boltzmann weight of an associated classical lattice gas problem defined on the same spatial lattice, which has the same density correlations as the ground state of the quantum problem. The kinetic terms associated with the pair of clusters $c$ and $c^{\prime}$ is
@@
the Hamiltonian can be recast into the form  [^1] This form is similar to that of the interacting boson models in the continuum studied in Ref. Kane et al. (1991) where the engineered Hamiltonian can be written into the form $\hat{\mathcal{H}}=$ $\hat{P}^{-1}\left(\nabla\psi^{\dagger}\right)\hat{P}\cdot\hat{P}\left(\nabla\psi\right)\hat{P}^{-1}.$ (8) with $\hat{P}$ related to a corresponding classical Boltzmann weight determined by the boson configuration.
$\hat{H}_{c,c^{\prime}}=$ $\hat{P}^{-1}\left(\hat{B}^{\dagger}_{c}-\hat{B}^{\dagger}_{c^{\prime}}\right)\hat{P}\hat{P}_{c}\hat{P}_{c^{\prime}}\hat{P}\left(\hat{B}_{c}-\hat{B}_{c^{\prime}}\right)\hat{P}^{-1}.$ (9)
Since these Hamiltonians can be written into the form of $\hat{O}^{\dagger}\hat{O}$ with $\hat{O}=\left(\hat{P}_{c}\hat{P}_{c^{\prime}}\right)^{1/2}\hat{P}\left(\hat{B}_{c}-\hat{B}_{c^{\prime}}\right)\hat{P}^{-1}$, they are always positive semi-definite.
@@
$\left(\hat{B}_{c}-\hat{B}_{c^{\prime}}\right)\left(\hat{b}^{\dagger}_{0}\right)^{M}|0\rangle=0$ (10)
@@
is the ground-state of $\hat{H}$, since $\hat{H}$ is non-negative and $\hat{H}|M\rangle=0$. As long as $\tau_{c,c^{\prime}}$ is generic enough to connect all the configurations of $\{n\}$, the ground state is unique (up to potential degeneracies associated with spontaneously broken symmetries or topological order) due to the Perron-Frobenius theorem.
@@
Dynamical density fluctuations: Besides the ground states, the dynamics of the density fluctuations can also be inferred. For concreteness, in this and the next section, we will consider the case of translationally invariant systems on regular lattices so that the excitations can be assigned a Bloch wave vector, $\bm{q}$.
@@
$f(\bm{q})\equiv$ $\frac{1}{\mathsf{V}}\left\langle\alpha\left|\hat{\rho}_{\bm{q}}^{\dagger}[\hat{H},\hat{\rho}_{\bm{q}}]\right|\alpha\right\rangle$ (21)
@@
Quantum dimer models : Gutzwiller projection is not the only case in which the present approach can be used to study model problems defined in constrained Hilbert places. Here, we show how this scheme can also unify quantum dimer models (QDM). In this case, the sites are taken to lie on the edges of a given lattice (forming its generalized ‘Lieb lattice’), and the bosons are considered ‘dimers.’ The dimer covering constraint can be achieved by
@@
where $N_{\mathsf{v}}\equiv\sum_{i\in\partial\mathsf{v}}n_{i}$ is the number of dimers occupying the edges connected to the vertex $\mathsf{v}$ (represented by $i\in\partial\mathsf{v}$), and $\bar{N}=\bar{n}\mathsf{C}/2$ is the average number of $N_{\mathsf{v}}$ ($\mathsf{C}$ is the coordination number of the lattice). For $10<\bar{n}<$ 1, the corresponding $\hat{P}$ projects onto states with at most one dimer on each edge as $u\rightarrow\infty$, and either $\lfloor\bar{N}\rfloor$ or $\lfloor\bar{N}\rfloor+1$ dimers touching each vertex as $u^{\prime}\to\infty$. Specifically, for $\bar{N}=1$, a hard-core dimer constraint is enforced, and hence the reference classical statistical mechanics model is the much-studied hard-core dimer model. On the other hand, for $\bar{N}=2$, the corresponding classical model sums over non-intersecting, undirected, fully-packed loops, another famous model in classical statistical mechanics  [^5] More generally, for any $\bar{N}\leq 2$, the reference classical statistical mechanics model corresponding to undirected, non-intersecting loops at general packing can be obtained by substituting $\left(N_{\mathsf{v}}-\bar{N}\right)^{2}\to N_{\mathsf{v}}\left(N_{\mathsf{v}}-2\right)^{2}$ in Eq. LABEL:loops, and then taking the familiar limit $u$ and $u^{\prime}\to\infty$..
@@
Other models with constrained Hilbert spaces: Further extending this methodology, one can also obtain ad hoc boson representations of certain Levin-Wen string-net models Levin and Wen (2005); Lin et al. (2021), Kitaev’s toric code Kitaev (2006) or quantum vertex models Balasubramanian et al. (2022b); Ardonne et al. (2004); Hermele et al. (2004); Balasubramanian et al. (2022a). The strategy is to introduce several flavors of bosons on the lattice edge to represent the possible ‘string types’ and then use an appropriate $K(\{n\})$ to impose a single-occupancy constraint on each edge and a constraint at each vertex so that surrounding string types obey the ‘branching rules.’ By considering the quantum resonance of the strings consistent with the rules, one can, in this way, naturally obtain the solvable Hamiltonians describing certain deconfined gauge theories.
@@
The addition of a density interaction $w(\delta n)^{2}/2$ to the effective action makes the system unstable to phase separation (as usual) for attractive $w<0$. In contrast, for repulsive $w>0$, it keeps the gapless manifold $\mathcal{M}$ but changes the boson excitation dispersion near $\mathcal{M}$ from quadratic to linear: $E(\bm{q})\sim\sqrt{wf(\bm{q})}$, which results in a spatially decaying boson Green’s function $\propto\langle\mathrm{e}^{\mathrm{i}[\theta(\bm{r})-\theta(\bm{0})]}\rangle$ with a similar character as for a Fermi liquid with the same manifold $\mathcal{M}$ of gapless modes (e.g. a surface in 3D or a nodal line in 2D)  [^6] It should be noted that the state constructed here is different from the Bose liquid studied in Ref. Motrunich and Fisher (2007) in many aspects, including its specific heat and the relation between the particle density and the volume enclosed by the Bose surfaces. .
Discussion: In this work, we have developed a unified approach for constructing “frustration-free” Hamiltonians for number-conserving bosons whose exact ground-state properties can be inferred from the properties of an associated classical statistical mechanical problem on the same lattice. For the known solvable models, our scheme provides not only an alternative representation that might be realizable on engineered platforms like optical lattices but also a natural route from a strongly constrained problem to a free (trivial) limit of the boson problem. When the constrained limit of the problem exhibits an exotic quantum phase of matter, this evolution must drive exotic transitions out of them, whose properties may be studied with these engineered models. The critical points constructed this way are “conformal quantum critical points” Ardonne et al. (2004) with unusual dynamical properties Castelnovo et al. (2005); Isakov et al. (2011).
@@
Acknowledgements: We thank Jonah Herzog-Arbeitman, Steven Girvin, Vladimir Calvera, Kyung-Su Kim, and Eduardo Fradkin for helpful communications. This work was supported, in part, by the U. S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, under Award No. DE-SC0020045.